# Federated Learning With Energy Harvesting Devices: An MDP Framework

### v1

@@ Abstract
Federated learning (FL) requires edge devices to perform local training and exchange information with a parameter server, leading to substantial energy consumption. A critical challenge in practical FL systems is the rapid energy depletion of battery-limited edge devices, which curtails their operational lifespan and affects the learning performance. To address this issue, we apply energy harvesting technique in FL systems to extract ambient energy for continuously powering edge devices. We first establish the convergence bound for the wireless FL system with energy harvesting devices, illustrating that the convergence is impacted by partial device participation and packet drops, both of which depend on the energy supply. To accelerate the convergence, we formulate a joint device scheduling and power control problem and model it as a Markov decision process (MDP). By solving this MDP, we derive the optimal transmission policy and demonstrate that it possesses a monotone structure with respect to the battery and channel states. To overcome the curse of dimensionality caused by the exponential complexity of computing the optimal policy, we propose a low-complexity algorithm, which is asymptotically optimal as the number of devices increases. Furthermore, for unknown channels and harvested energy statistics, we develop a structure-enhanced deep reinforcement learning algorithm that leverages the monotone structure of the optimal policy to improve the training performance. Finally, extensive numerical experiments on real-world datasets are presented to validate the theoretical results and corroborate the effectiveness of the proposed algorithms.
@@
Federated learning, energy harvesting, Markov decision process, monotone structure, the curse of dimensionality, deep reinforcement learning.
@@ I Introduction
Machine learning is a cornerstone technology driving artificial intelligence (AI) in diverse fields, such as autonomous driving [1] and the Internet of Things [2]. Traditional centralized machine learning requires the aggregation and processing of distributed data on a central server, which leads to increased data traffic and raises concerns about data privacy and security. To address these issues, federated learning (FL) has emerged as a novel distributed machine learning paradigm that enables collaborative model training without uploading raw data to a central server[3]. In FL, edge devices (e.g., mobile phones and sensors) conduct local training on their respective datasets and subsequently exchange information (e.g., model parameters or gradients) with a parameter server. Due to its distinctive architecture, FL can alleviate the computation and communication burdens and preserve the data privacy of involved edge devices [4, 5].
@@
In practical scenarios, the performance of FL is severely affected by radio resources, such as transmission power and bandwidth. The limited radio resources among participating edge devices hinder data transmission and increase communication latency, degrading the learning performance. To tackle these issues, numerous papers have investigated communication resource management for wireless FL systems. In [6], the authors investigated a joint radio resource allocation and user selection problem for wireless FL, aiming to minimize the training loss under delay and energy consumption constraints. A similar resource allocation problem of achieving the optimal tradeoff between convergence speed and power consumption for wireless FL was studied in [7]. In [8], the authors developed a joint computation and transmission scheme for wireless FL under a more challenging setting with heterogeneous data. While the aforementioned studies explored static resource allocation according to the statistical properties of wireless channels, other research focused on the dynamic allocation of communication resources. The authors of [9] proposed a dynamic power control policy based on instant channel state information (CSI) to speed up the convergence of over-the-air FL. Similarly, a dynamic device scheduling and power control scheme for over-the-air FL was developed in [10]. Bandwidth allocation for wireless FL has also been studied extensively. In [11], the authors investigated a joint device selection and bandwidth allocation problem for wireless FL, devising an online algorithm to maximize the number of participating devices in each round. A similar problem incorporating latency constraints was studied in [12]. Additionally, the authors of [13] studied a joint power and bandwidth allocation problem to simultaneously minimize the power consumption, computation cost, and time cost of wireless FL systems.
@@
Despite the extensive research, all the aforementioned studies have either considered finite energy supply for edge devices or completely neglected the energy limitations. However, merely utilizing the finite energy stored in the batteries of edge devices will lead to a reduced operational lifespan. Once the battery is depleted, edge devices cannot contribute to the training process of FL, which degrades the learning performance. Additionally, frequent battery replacements and recharges are impractical, especially for devices deployed in remote or inaccessible environments. To address this issue, energy harvesting (EH) FL has emerged as a promising solution. This technique collects ambient energy (e.g., solar, wind or thermoelectric energy) to continuously power edge devices, thereby extending their operational lifetime and improving the learning performance. Furthermore, EH FL uses clean and renewable energy sources, making it more environmentally friendly compared to conventional FL with battery-powered edge devices. Several recent studies have investigated EH FL. In [14], a client scheduling algorithm was introduced to accelerate the convergence of FL over EH-enabled wireless networks. A joint device selection and resource allocation problem was investigated in [15] to reduce training latency and improve the data utility of EH FL. The authors of [16] extended this problem to a more sophisticated system, where the server employs massive multiple-input multiple-output (MIMO) to serve clients. Exploiting the over-the-air computation, the authors of [17] proposed a bandwidth-efficient scheme for joint client selection and energy management in EH FL. In [18], a joint client selection and bandwidth allocation scheme was designed for FL with EH devices, aiming to minimize the ages of local models and the latency of global iterations. To accelerate the convergence of EH FL, the authors of [19] formulated a joint transceiver design and device scheduling problem and developed a Lyapunov-based online optimization algorithm to solve it.
@@
The prior works [14, 15, 16, 17, 18] on EH FL concentrate on designing transmission strategies for one typical time slot. However, the overall performance of FL is affected by the cumulative model aggregation errors over the entire training period, necessitating careful utilization of harvested energy from a long-term perspective. We note that [19] is an early attempt to optimize the long-term efficiency of EH FL. However, its online optimization algorithm fails to exploit the statistical knowledge of system dynamics, e.g., the distributions of the channel gains and the amount of harvested energy, resulting in suboptimal learning performance. Thus, in this paper, we are motivated to develop a Markov decision process (MDP) framework for designing the optimal transmission policy of EH FL. The proposed MDP framework utilizes the statistical knowledge of channels and harvested energy to address their inherent stochastic nature, thereby enabling efficient utilization of harvested energy and enhanced learning performance. The main contributions are summarized as follows.
@@ II-A Federated Learning Model
We consider a wireless FL system, where $N$ edge devices collaboratively train a common learning model $\boldsymbol{w}=[w_{1},\ldots,w_{d}]\in\mathbb{R}^{d}$ under the coordination of a parameter server, as shown in Fig. 1. Each edge device $n$ has a local dataset $\mathcal{D}_{n}=\{(\boldsymbol{x}_{n,i},y_{n,i})\}_{i=1}^{D_{n}}$ with $D_{n}=|\mathcal{D}_{n}|$ data samples, where $\boldsymbol{x}_{n,i}\in\mathbb{R}^{d}$ is the $i$-th input feature vector of device $n$ and $y_{n,i}$ is its corresponding label. Then, the global dataset is given by $\mathcal{D}=\bigcup_{n=1}^{N}\mathcal{D}_{n}$ with $D=\sum_{n=1}^{N}D_{n}$ data samples. Let $f(\boldsymbol{w};\boldsymbol{x}_{n,i},y_{n,i})$ denote the sample-wise loss function, which quantifies the loss of learning model $\boldsymbol{w}$ for data sample $(\boldsymbol{x}_{n,i},y_{n,i})$. The goal of the FL system is to train an optimal global model $\boldsymbol{w}^{*}$ that minimizes the global loss function $F(\boldsymbol{w})$, i.e.,
$\min_{\boldsymbol{w}}\;F(\boldsymbol{w})=\sum_{n=1}^{N}\frac{D_{n}}{D}F_{n}(\boldsymbol{w}),$ (1)
where $F_{n}(\boldsymbol{w})$ is the local loss function of device $n$, defined as
$F_{n}(\boldsymbol{w})$ $=\frac{1}{D_{n}}\!\sum_{(\boldsymbol{x}_{n,i},y_{n,i})\in\mathcal{D}_{n}}\!\!f(\boldsymbol{w};\boldsymbol{x}_{n,i},y_{n,i}).$ (2)
To solve the learning problem (1) in a distributed manner, the FL system conducts the following three steps iteratively under the coordination of the server:
  • The server schedules a subset of edge devices and broadcasts the global model to them.
  • Each scheduled device computes its local gradient on its own local dataset by using the received learning model, and then uploads the local gradient to the server.
  • The server aggregates local gradients of scheduled devices and then updates the global model.
Each iteration consisting of these three steps is called a time slot, indexed by $t$. The whole training process continues until reaching a preset maxmimum number of time slots $T$.
@@ II-B Local Computation Model
At time slot $t$, the server first schedules a subset of edge devices $\Omega_{t}\subseteq[N]$ and then broadcasts the current global model $\boldsymbol{w}_{t}$ to these selected devices. We assume that the server can broadcast the global model to edge devices without any error, due to its strong communication capability (e.g., large transmission power). Let $\beta_{n,t}\in\left\{0,1\right\}$ denote the scheduling indicator, with $\beta_{n,t}=1$ if device $n$ is scheduled at time slot $t$, and $\beta_{n,t}=0$ otherwise. Then, we have $\Omega_{t}=\left\{n\,|\,\beta_{n,t}=1,n\in[N]\right\}$. After receiving the global model $\boldsymbol{w}_{t}$, each scheduled device $n\in\Omega_{t}$ computes its local gradient ${\boldsymbol{g}}_{n,t}$ on its own local dataset as follows:
${\boldsymbol{g}}_{n,t}=\frac{1}{D_{n}}\!\sum_{(\boldsymbol{x}_{n,i},y_{n,i})\in{\mathcal{D}}_{n}}\!\!\nabla f(\boldsymbol{w}_{t};\boldsymbol{x}_{n,i},y_{n,i}).$ (3)
We consider a single-iteration gradient update in this work, and it can be extended to scenarios with multiple local iterations. The energy consumption for local computation of device $n$ at time slot $t$ is given by [20]
$e_{n,t}^{\text{cmp}}=I_{n}^{2}c_{n}D_{n},$ (4)
where $I_{n}$ is the CPU cycle frequency of device $n$, and $c_{n}$ is the number of CPU cycles required for device $n$ to process one data sample. Since the local computation energy $e_{n,t}$ is identical across time slots, we omit its subscript $t$ henceforth.
@@ II-C Communication Model
After local computation, the scheduled edge devices send their local gradients to the server for global aggregation. For the uplink transmission, we adopt an orthogonal frequency division multiple access (OFDMA) technique, where each device occupies one resource block to upload its local gradient. We consider a non-trivial scenario where no more than $R<N$ devices can upload their local gradients at one time slot due to the limited bandwidth constraint. Thus, we have
@@
where $s_{n}$ denotes the size of the $n$-th device’s local gradient. The total energy consumption of device $n$ at time slot $t$ is given by
$e_{n,t}=e_{n}^{\text{cmp}}+e_{n,t}^{\text{com}}.$ (8)
In this paper, we assume that the local gradient of each device is transmitted as a single packet in the uplink. Due to the limited bandwidth and transmission power of edge devices, the packet error rate of uplink transmission cannot be ignored. Specifically, the packet error rate of the uplink transmission from device $n$ to the server at time slot $t$ is given by [21]
@@
where $\sigma$ is a constant parameter, namely, the waterfall threshold [21]. Let $\zeta_{n,t}\in\left\{0,1\right\}$ denote the indicator of successful transmission, where $\zeta_{n,t}=0$ indicates that the received local gradient of device $n$ at time slot $t$ is dropped due to the occurrence of packet error, and $\zeta_{n,t}=1$ otherwise. Then, we have
@@
In the considered wireless FL system, the server employs successfully received local gradients from the scheduled devices to update the global model as follows:
$\boldsymbol{w}_{t+1}=\boldsymbol{w}_{t}-\eta\frac{\sum_{n=1}^{N}\beta_{n,t}\zeta_{n,t}{\boldsymbol{g}}_{n,t}}{\sum_{n=1}^{N}\beta_{n,t}\zeta_{n,t}},$ (11)
where $\eta$ is the learning rate. If all devices lack sufficient energy to upload local gradients to the server, or if all transmitted local gradients are lost, the server can utilize the previous global model, i.e., $\boldsymbol{w}_{t+1}=\boldsymbol{w}_{t}$.
@@ III-A Convergence Analysis
To facilitate the convergence analysis, we make the following standard assumptions, which are widely adopted in the literature on FL [8, 22, 23, 24, 25, 26, 27].
@@
(Smoothness) The local loss function $F_{n}(\boldsymbol{w})$ is $L$-smooth. For any $\boldsymbol{a},\boldsymbol{b}\in\mathbb{R}^{d}$, $n\in[N]$, we have
@@
(Strong Convexity) The local loss function $F_{n}(\boldsymbol{w})$ is $\mu$-strongly convex. For any $\boldsymbol{a},\boldsymbol{b}\in\mathbb{R}^{d}$, $n\in[N]$, we have
$\begin{split}F_{n}(\boldsymbol{a})-F_{n}(\boldsymbol{b})\geq\nabla F_{n}(\boldsymbol{b})^{\mathsf{T}}(\boldsymbol{a}-\boldsymbol{b})+\frac{\mu}{2}\|\boldsymbol{a}-\boldsymbol{b}\|^{2}.\end{split}$ (15)
@@
(Bounded Gradient) For any model $\boldsymbol{w}$ and data sample $(\boldsymbol{x}_{n,i},y_{n,i})$, there exists a constant $\rho>0$ satisfying
$\left\|\nabla f(\boldsymbol{w};\boldsymbol{x}_{n,i},y_{n,i})\right\|^{2}\leq\rho.$ (16)
Based on the assumptions above, we characterize the upper bound of the expected optimality gap in the following theorem.
@@
$\mathbb{E}\left[F(\boldsymbol{w}_{T+1})-F(\boldsymbol{w}^{*})\right]$ (17)
$\leq$ $A^{T}\left(F(\boldsymbol{w}_{1})-F(\boldsymbol{w}^{*})\right)$
$+\frac{\eta\rho}{2N}\sum_{t=1}^{T}A^{T-t}\sum_{n=1}^{N}\frac{1-\beta_{n,t}(1-q_{n,t}(h_{n,t},p_{n,t}))}{D_{n}},$
where $A=1-\eta\mu\in(0,1)$ and the learning rate satisfies $\eta\leq\min\left\{\frac{1}{L},1\right\}$.
@@
The first term on the right-hand-side (RHS) of (17) denotes the initial error caused by the gap between the initial model $\boldsymbol{w}_{1}$ and the optimal model $\boldsymbol{w}^{*}$, while the second term corresponds to the error caused by partial device participation and packet drops. Specifically, the convergence rate of wireless FL can be improved by scheduling more devices and increasing their transmission power. However, in practical scenarios, both device scheduling and transmission power are constrained by the battery energy level. Thus, the transmission policy should be carefully designed to utilize the limited and uncertain harvested energy effectively.

### v2

@@ Abstract
Federated learning (FL) necessitates that edge devices conduct local training and communicate with a parameter server, resulting in significant energy consumption. A key challenge in practical FL systems is the rapid depletion of battery-limited edge devices, which limits their operational lifespan and impacts learning performance. To tackle this issue, we implement energy harvesting techniques in FL systems to capture ambient energy, thereby providing continuous power to edge devices. We first establish the convergence bound for the wireless FL system with energy harvesting devices, illustrating that the convergence is affected by partial device participation and packet drops, both of which depend on the energy supply. To accelerate the convergence, we formulate a joint device scheduling and power control problem and model it as a Markov decision process (MDP). By solving this MDP, we derive the optimal transmission policy and demonstrate that it possesses a monotone structure with respect to the battery and channel states. To overcome the curse of dimensionality caused by the exponential complexity of computing the optimal policy, we propose a low-complexity algorithm, which is asymptotically optimal as the number of devices increases. Furthermore, for unknown channels and harvested energy statistics, we develop a structure-enhanced deep reinforcement learning algorithm that leverages the monotone structure of the optimal policy to improve the training performance. Finally, extensive numerical experiments on real-world datasets are presented to validate the theoretical results and corroborate the effectiveness of the proposed algorithms.
@@
Federated learning, energy harvesting, Markov decision process, monotone structure, curse of dimensionality, deep reinforcement learning.
@@ I Introduction
Machine learning is a cornerstone technology driving artificial intelligence (AI) in diverse fields, such as autonomous driving [1] and the Internet of Things [2]. Traditional centralized machine learning requires the aggregation and processing of distributed data on a central server, which leads to increased data traffic and raises concerns about data privacy and security. To address these issues, federated learning (FL) has emerged as a novel distributed machine learning paradigm that enables collaborative model training without uploading raw data to a central server[3, 4]. In FL, edge devices (e.g., mobile phones and sensors) conduct local training on their respective datasets and subsequently exchange information (e.g., model parameters or gradients) with a parameter server. Due to its distinctive architecture, FL can alleviate the computation and communication burdens and preserve the data privacy of involved edge devices [5, 6].
@@
In practical scenarios, the performance of FL is severely affected by radio resources, such as transmission power and bandwidth. The limited radio resources among participating edge devices hinder data transmission and increase communication latency, thereby degrading the learning performance. To tackle these issues, numerous papers have investigated communication resource management for wireless FL systems. In [7], the authors investigated a joint radio resource allocation and user selection problem for wireless FL, aiming to minimize the training loss under delay and energy consumption constraints. A similar resource allocation problem of achieving the optimal tradeoff between convergence speed and power consumption for wireless FL was studied in [8]. In [9], the authors developed a joint computation and transmission scheme for wireless FL under a more challenging setting with heterogeneous data. While the aforementioned studies explored static resource allocation according to the statistical properties of wireless channels, other research focused on the dynamic allocation of communication resources. The authors of [10] proposed a dynamic power control policy based on instant channel state information (CSI) to speed up the convergence of over-the-air FL. Similarly, a dynamic device scheduling and power control scheme for over-the-air FL was developed in [11]. Bandwidth allocation for wireless FL has also been studied extensively. In [12], the authors investigated a joint device selection and bandwidth allocation problem for wireless FL, devising an online algorithm to maximize the number of participating devices in each round. A similar problem incorporating latency constraints was studied in [13]. Additionally, the authors of [14] studied a joint power and bandwidth allocation problem to simultaneously minimize the power consumption, computation cost, and time cost of wireless FL systems.
@@
Despite the extensive research, all the aforementioned studies have either considered finite energy supply for edge devices or completely neglected the energy limitations. However, merely utilizing the finite energy stored in the batteries of edge devices will lead to a reduced operational lifespan. Once the battery is depleted, edge devices cannot contribute to the training process of FL, which degrades the learning performance. Additionally, frequent battery replacements and recharges are impractical, especially for devices deployed in remote or inaccessible environments. To address this issue, energy harvesting (EH) FL has emerged as a promising solution. This technique collects ambient energy (e.g., solar, wind or thermoelectric energy) to continuously power edge devices, thereby extending their operational lifetime and improving the learning performance. Furthermore, EH FL uses clean and renewable energy sources, making it more environmentally friendly compared to conventional FL with battery-powered edge devices. Several recent studies have investigated EH FL. In [15], a distributed stochastic gradient descent (SGD) algorithm was developed under intermittent and heterogeneous energy arrivals, with provable convergence guarantees. The subsequent work [16] extended this algorithm by proposing a scalable client scheduling and model update scheme based on each client’s energy renewal cycle, achieving notable improvements over energy-agnostic baselines. In [17], a client scheduling algorithm was introduced to accelerate the convergence of FL over EH-enabled wireless networks. A joint device selection and resource allocation problem was investigated in [18] to reduce training latency and improve the data utility of EH FL. The authors of [19] extended this problem to a more sophisticated system, where the server employs massive multiple-input multiple-output (MIMO) to serve clients. Exploiting the over-the-air computation technique, bandwidth-efficient transmission schemes tailored for EH FL were proposed in [20, 21]. In [22], a joint client selection and bandwidth allocation scheme was designed for FL with EH devices, aiming to minimize the ages of local models and the latency of global iterations. To accelerate the convergence of EH FL, the authors of [23] formulated a joint transceiver design and device scheduling problem and developed a Lyapunov-based online optimization algorithm to solve it.
@@
The prior works [15, 16, 17, 18, 19, 20, 21, 22] on EH FL concentrate on designing transmission strategies for one typical time slot. However, the overall performance of FL is affected by the cumulative model aggregation errors over the entire training period, necessitating careful utilization of harvested energy from a long-term perspective. We note that [23] is an early attempt to optimize the long-term efficiency of EH FL. However, its online optimization algorithm fails to exploit the statistical knowledge of system dynamics, e.g., the distributions of the channel gains and the amount of harvested energy, resulting in suboptimal learning performance. Thus, in this paper, we are motivated to develop a Markov decision process (MDP) framework for designing the optimal transmission policy of EH FL. The proposed MDP framework utilizes the statistical knowledge of channels and harvested energy to address their inherent stochastic nature, thereby enabling efficient utilization of harvested energy and enhanced learning performance. The main contributions are summarized as follows.
@@ II-A Federated Learning Model
We consider a wireless FL system, where $N$ edge devices collaboratively train a common learning model $\boldsymbol{w}=[w_{1},\ldots,w_{d}]\in\mathbb{R}^{d}$ under the coordination of a parameter server, as shown in Fig. 1. Each device $n$ is associated with a local loss function $F_{n}(\boldsymbol{w})=\mathbb{E}_{\xi_{n}\sim\mathcal{D}_{n}}\left[f_{n}(\boldsymbol{w};\xi_{n})\right]$, where $f_{n}(\boldsymbol{w};\xi_{n})$ denotes the sample-wise loss function that quantifies the loss of learning model $\boldsymbol{w}$ on the data sample $\xi_{n}$, and $\mathcal{D}_{n}$ is the local dataset of device $n$. The goal of the FL system is to train an optimal global model $\boldsymbol{w}^{*}$ that minimizes the global loss function $F(\boldsymbol{w})$, i.e.,
$\min_{\boldsymbol{w}}\;F(\boldsymbol{w})=\frac{1}{N}\sum_{n=1}^{N}F_{n}(\boldsymbol{w}).$ (1)
Problem (1) is known as empirical risk minimization, which arises in many ML problems. To solve the learning problem (1) in a distributed manner, the FL system conducts the following three steps iteratively under the coordination of the server:
  • The server schedules a subset of edge devices and broadcasts the current global model to them.
  • Each scheduled device performs multiple steps of local SGD on its private dataset, and then uploads its local model update to the server.
  • The server aggregates the local updates received from the participating devices and then refines the global model.
Each iteration consisting of these three steps is called a time slot, indexed by $t$. The whole training process continues until reaching a preset maximum number of time slots $T$.
@@ II-B Local Computation Model
At time slot $t$, the server first schedules a subset of edge devices $\Omega_{t}\subseteq[N]$ and then broadcasts the current global model $\boldsymbol{w}_{t}$ to these selected devices. We assume that the server can broadcast the global model to edge devices without any error, due to its strong communication capability (e.g., large transmission power). Let $\beta_{n,t}\in\left\{0,1\right\}$ denote the scheduling indicator, with $\beta_{n,t}=1$ if device $n$ is scheduled at time slot $t$, and $\beta_{n,t}=0$ otherwise. Then, we have $\Omega_{t}=\left\{n\,|\,\beta_{n,t}=1,n\in[N]\right\}$. After receiving the global model $\boldsymbol{w}_{t}$, each scheduled device $n\in\Omega_{t}$ initialize the model for local SGD by setting $\boldsymbol{w}_{n,t}^{(1)}=\boldsymbol{w}_{t}$. Then, for each local iteration $k\in\{1,2,\ldots,K\}$, device $n$ samples a mini-batch $\hskip 2.0pt\boldsymbol{\xi}_{n,t}^{(k)}\hskip 0.5pt\subseteq\hskip 0.5pt\mathcal{D}_{i}$ from its local dataset and computes the stochastic gradient as follows:
$\boldsymbol{g}_{n,t}^{(k)}=\nabla f_{n}(\boldsymbol{w}_{n,t}^{(k)};\boldsymbol{\xi}_{n,t}^{(k)}).$ (2)
The local model is then updated as follows:
$\boldsymbol{w}_{n,t}^{(k+1)}=\boldsymbol{w}_{n,t}^{(k)}-\eta\,\boldsymbol{g}_{n,t}^{(k)},$ (3)
where $\eta>0$ denotes the learning rate. The local computation energy consumption for device $n$ at time slot $t$ can be expressed as [24]
$e_{n,t}^{\text{cmp}}=\beta_{n,t}KI_{n}^{2}c_{n}|\boldsymbol{\xi}_{n}|,$ (4)
where $I_{n}$ is the CPU cycle frequency of device $n$, $c_{n}$ denotes the number of CPU cycles required to process one data sample, and $|\boldsymbol{\xi}_{n}|$ is the (constant) mini-batch size used in each local iteration.
@@ II-C Communication Model
After local computation, each scheduled device $n$ send its local model update $\Delta\boldsymbol{w}_{n,t}=\boldsymbol{w}^{(K+1)}_{n,t}-\boldsymbol{w}_{t}$ to the server for global aggregation. For the uplink transmission, we adopt an orthogonal frequency division multiple access (OFDMA) technique, where each device occupies one resource block to upload its local update. We consider a non-trivial scenario where no more than $R<N$ devices can upload their local updates at one time slot due to the limited bandwidth constraint. Thus, we have
@@
where $s_{n}$ denotes the size of the $n$-th device’s local model update. The total energy consumption of device $n$ at time slot $t$ is given by
$e_{n,t}=e_{n,t}^{\text{cmp}}+e_{n,t}^{\text{com}}.$ (8)
In this paper, we assume that the local model update of each device is transmitted as a single packet in the uplink. Due to the limited bandwidth and transmission power of edge devices, the packet error rate of uplink transmission cannot be ignored. Specifically, the packet error rate of the uplink transmission from device $n$ to the server at time slot $t$ is given by [25]
@@
where $\sigma$ is a constant parameter, namely, the waterfall threshold [25]. Let $\zeta_{n,t}\in\left\{0,1\right\}$ denote the indicator of successful transmission, where $\zeta_{n,t}=0$ indicates that the received local update of device $n$ at time slot $t$ is dropped due to the occurrence of packet error, and $\zeta_{n,t}=1$ otherwise. Then, we have
@@
In the considered wireless FL system, the server employs successfully received local updates from the scheduled devices to refine the global model as follows:
$\boldsymbol{w}_{t+1}=\boldsymbol{w}_{t}+\frac{\sum_{n=1}^{N}\beta_{n,t}\zeta_{n,t}\Delta\boldsymbol{w}_{n,t}}{\sum_{n=1}^{N}\beta_{n,t}\zeta_{n,t}}$ (11)
If all devices lack sufficient energy to upload local gradients to the server, or if all transmitted local gradients are lost, the server can utilize the previous global model, i.e., $\boldsymbol{w}_{t+1}=\boldsymbol{w}_{t}$.
@@ III-A Convergence Analysis
To facilitate the convergence analysis, we make the following standard assumptions, which are widely adopted in the literature on FL [9, 26, 27, 28, 29].
@@
(Convexity and Smoothness) The local loss function $F_{n}(\boldsymbol{w})$ is convex and $L$-smooth. For any $\boldsymbol{a},\boldsymbol{b}\in\mathbb{R}^{d}$, $n\in[N]$, we have
@@
(Bounded Variance) The stochastic gradient of each device is unbiased and has a bounded local variance, i.e., for any $\boldsymbol{w}\in\mathbb{R}^{d}$, $i\in[m]$, we have
$\mathbb{E}_{\mathcal{\xi}\sim\mathcal{D}_{i}}\left[\left\|\nabla f_{i}(\boldsymbol{w},\mathcal{\xi})-\nabla F_{i}(\boldsymbol{w})\right\|^{2}\right]\leq\sigma_{l}^{2}.$ (15)
The local loss functions have a dissimilarity bound that characterizes their heterogeneity. For any $\boldsymbol{w}\in\mathbb{R}^{d}$, $i\in[m]$, we have
$\left\|\nabla F_{i}(\boldsymbol{w})-\nabla F(\boldsymbol{w})\right\|^{2}\leq\sigma_{g}^{2}.$ (16)
@@
(Bounded Stochastic Gradient) The expected squared norm of the stochastic gradients is uniformly bounded, i.e., for any $\boldsymbol{w}\in\mathbb{R}^{d}$, $i\in[m]$, we have
$\mathbb{E}\left[\left\|\nabla f_{i}(\boldsymbol{w},\mathcal{\xi})\right\|^{2}\right]\leq G^{2}.$ (17)
Based on the assumptions above, we characterize the convergence bound of wireless FL with EH for general convex loss functions. Specifically, we analyze the convergence of the average model, which is defined as
$\boldsymbol{\bar{w}}_{t}^{(k)}=\frac{1}{N}\sum_{n=1}^{N}\boldsymbol{w}_{n,t}^{(k)}.$ (18)
@@
$\mathbb{E}\Biggl{[}\frac{1}{KT}\sum_{t=1}^{T}\sum_{k=1}^{K}\bigl{(}F(\bar{\boldsymbol{w}}_{t}^{(k)})-F(\boldsymbol{w}^{\star})\bigr{)}\Biggr{]}$ (19)
$\leq$ $\frac{2L\|\boldsymbol{w}_{1}-\boldsymbol{w}^{*}\|^{2}}{\sqrt{KT}}+\frac{\sigma_{l}^{2}}{4LN\sqrt{KT}}+\frac{\sigma_{l}^{2}}{4LT}+\frac{9\sigma_{g}^{2}K}{8LT}$
$+\frac{2LG^{2}}{N\sqrt{KT}}\sum_{t=1}^{T}\sum_{n=1}^{N}\left(1-\beta_{n,t}(1-q_{n,t}(h_{n,t},p_{n,t}))\right),$
where the learning rate satisfies $\eta=\frac{1}{4L\sqrt{KT}}$.
@@
The first four terms on the right-hand side (RHS) of (LABEL:convergence_bound) arise from the initial error (i.e., the gap between the initial model ${\boldsymbol{w}}_{1}$ and the optimal model $\boldsymbol{w}^{*}$) as well as other learning-related factors (e.g., device heterogeneity and data heterogeneity). These terms gradually diminish as the number of time slots $T$ increases. In contrast, the final term represents the persistent error introduced by partial device participation and transmission packet drops, and this error does not vanish as $T$ grows. In particular, the convergence rate of wireless FL can be improved by scheduling more devices and increasing their transmission power. However, in practical scenarios, both device scheduling and transmission power are constrained by the battery energy level. Thus, the transmission policy should be carefully designed to utilize the limited and uncertain harvested energy effectively.